for a first order reaction t_0.75 is 138.6 seconds.its specific reaction rate is

Grade:12

1 Answers

Sunil Kumar FP
10 years ago
t,75=2t.5
t1/2=1/2*t.75
1/2*138.6=69.3 sec
k-.693/t(1/2)
t(1/2)=.01

thanks and regards
sunil kr
askIITian faculty